Part A Calculate the pH in 1.26 M CH3CO2H.

Part B

Calculate the concentrations of all species present (H3O+, CH3CO2−, CH3CO2H, and OH−) in 1.26 M CH3CO2H.

Express your answers using two significant figures separated by commas. Enter your answers in order listed in the question.

Part C

Calculate the pH in 9.50×10−2 M CH3CO2H.

Part D

Calculate the concentrations of all species present (H3O+, CH3CO2−, CH3CO2H, and OH−) in 9.50×10−2 M CH3CO2H.

Express your answers using two significant figures separated by commas. Enter your answers in order listed in the question.
Answer #1

a) pH of weak acid = 1/2(pka-logc)

                   = 1/2(4.756-log1.26)

                 = 2.3
[H3O+] = 10^(-PH)

     = 10^(-2.3) = 4.7x10-3 M

[H+] = [CH3COO-] = 4.7x10-3 M

CH3CO2H = 1.26 M

OH- = 10-14 / 4.7x10-3 M ==2.1 X 10-12 M

Post remaining separately .Or follow similar method so that you can learn